What is the mass (in grams) of 1.70M sodium sulfide solution that would react with 105 g of cadmium nitrate solution?


Na2S + Cd(NO3)2 = CdS+2NaNO3


The answer is 34.66g. I'd like to know what exactly are the equations that got to that?


1
Expert's answer
2022-03-21T21:29:03-0400

This question is a bit confusing as both masses are meant to be of pure substances, not solutions. The mass of the solution cannot be determined without knowing its density, and the result in that case would be different.


The molar masses are:

Cd(NO3)2 - 236.42 g/mol

Na2S - 78.04 g/mol

According to the equation, 1 mole of Cd(NO3)2 reacts with 1 mole os Na2S. Therefore, only one mass-to-mass conversion is required here:


"m(Na_2S)=105\ g\ Cd(NO_3)_2\times\frac{1\ mol\ Cd(NO_3)_2}{236.42\ g\ Cd(NO_3)_2}\times\frac{1\ mol\ Na_2S}{1\ mol\ Cd(NO_3)_2}\times\frac{78.04\ g\ Na_2S}{1\ mol\ Na_2S}=34.66\ g"


Answer: 34.66 g